A 9.5 V battery supplies a 3.5 mA current to a circuit for 7.0 h .

Part A: How much charge has been transferred from the negative to the positive terminal?

Part B: How much electric potential energy has been gained by the charges that passed through the battery?
Answer #1

I = Q/T

i.e ampere = coulomb / second

Now,

Part A: The amount of charge passed is:
Q = I * t = 3.5mA * 7h * 3600s/h = 88.2C

Part B: Electrical work is:
w = q * V = 88.2C * 9.5V = 837.9joule =838J